OpenStudy (anonymous):

I guess F is not a conservative vector field. \[F(x,y)=e^xsiny\hat{i}+e^xcosy\hat{j}\] \[\frac{\partial{P}}{\partial{y}}=\frac{\partial}{\partial{y}}|e^xsiny|=cosy\] \[\frac{\partial{Q}}{\partial{x}}=\frac{\partial}{\partial{x}}|e^xcosy|=e^x\] \[cosy\neq e^x\]

OpenStudy (anonymous):

The second one, I am not sure what you defined as which, but just to make sure. You are taking the partial derivative of e^x cosy ? with respect to x?

OpenStudy (anonymous):

\[\frac{\partial{P}}{\partial{y}}=\frac{\partial}{\partial{y}}|e^xsiny|=e^x cosy\]

OpenStudy (anonymous):

treat the respective values as constants for partial derivatives.

OpenStudy (anonymous):

exactly @91

OpenStudy (anonymous):

\[\frac{\partial{Q}}{\partial{x}}=\frac{\partial}{\partial{x}}|e^xcosy|=e^x cos y\] so there are conservative

OpenStudy (anonymous):

it is a conservative vector field.

OpenStudy (anonymous):

constant multplied to variable don't go away

OpenStudy (anonymous):

Oh ok. Thanks in the previous example that I did I got \[\frac{\partial}{\partial{y}}|2x-3y|=-3\] Which lead me to believe that the constant went away

OpenStudy (anonymous):

only if they are by themselve

OpenStudy (anonymous):

Makes sense. Thank you both!